# On hyperbolic links associated to Eulerian cycles on ideal right-angled hyperbolic $3$-polytopes

### v1

@@ Abstract.
We consider Eulerian cycles without transversal selfintersections in $4$-valent planar graphs. We prove that any cycle of this type in the graph of an ideal right-angled hyperbolic $3$-polytope corresponds to a hyperbolic link such that its complement consists of $4$-copies of this polytope glued according to its checkerboard coloring. Moreover, this link consists of trivially embedded circles bijectively corresponding to the vertices of the polytope. For such cycles we prove that the $3$-antiprism $A(3)$ (octahedron) has exactly $2$ combinatorially different cycles, the $4$-antiprism $A(4)$ has exactly $7$ combinatorially different cycles, and these cycles correspond to $7$ cycles (perhaps combinatorially equivalent) on any polytope different from antiprisms, and any antiprism $A(k)$ has at least $2$ combinatorially different cycles. The $2$-fold branched covering space corresponding to our link is a small cover over some simple $3$-polytope. This small cover is defined by a Hamiltonian cycle on it. We show that any compact hyperbolic $3$-polytope with a Hamiltonian cycle arises in this way.
@@ 1. Introduction
Our main result (Theorem 3.2) is that any Eulerian cycle without transversal selfintersections in the graph of any ideal right-angled $3$-polytope $P$ corresponds to a link whose circles bijectively correspond to vertices of $P$ and the complement is decomposed into $4$ copies of $P$. We show that the $2$-fold branched covering space corresponding to this link is a small cover build by the A.D. Mednykh [M90] construction from a Hamiltonian cycle on a simple $3$-polytope. We provide a method how to build such cycles on any ideal right-angled $3$-polytope.
@@ 2. Basic facts
If the components of $\partial Q\setminus\Gamma$ correspond to the pairs of colors $(1,2)$ and $(3,4)$, then $\tau=e_{1}+e_{2}=e_{3}+(e_{1}+e_{2}+e_{3})$ is the hyperelliptic involution on $N(Q,\widetilde{\Lambda}_{\Gamma})$. Moreover, the mapping $N(P,\widetilde{\Lambda}_{\Gamma})\to S^{3}$ is a $2$-fold branched covering with the following branch set (see details in [EE25, Section 4.5]). The edges of $Q$ not lying in $\Gamma$ form a perfect matching of the graph $G(Q)$ – a disjoint set of edges covering all the vertices. Then the preimage of this set in $N(Q,\widetilde{\Lambda}_{\Gamma})$ and in $S^{3}$ is a disjoint set of circles, each circle glued of two copies of the corresponding edge. This link is the brach set of the covering. The detailed description of this link is given in [G24]. In [VM99S2] this construction was generalized from Hamiltonian cycles to Hamiltonian theta-graphs and Hamiltonian $K_{4}$-subgraphs.
@@ 4. Edge-twists and the existence of nonselfcrossing Eulerian cycles
5. Links associated to Hamiltonian cycles on compact right-angled $3$-polytopes
@@
Every compact right-angled $3$-polytope $Q$ with a Hamiltonian cycle $\Gamma$ defines an ideal right-angled polytope $P$ obtained by shrinking to points the edges of the perfect matching consisting of the edges not in the cycle. The polytope $Q$ has an induced nonselfcrossing Eulerian cycle $\gamma$ such that $\Gamma=\Gamma_{\gamma}$. For the link $C_{\gamma}$ both the complement $S^{3}\setminus C_{\gamma}$ and the $2$-fold branched covering space have complete hyperbolic structures obtained by gluing right-angled polytopes.
@@
The Dodecahedron is a unique compact right-angled polytope with minimal number of facets (equal to $12$). Up to combinatorial symmetries it has a unique Hamiltonian cycle. In Fig. 7 we show this Hamiltonian cycle, and the way how the associated ideal right-angled polytope is obtained from $A(4)$ by a sequence of two restricted edge-twists. We also show another polytope corresponding to another Eulerian cycle on the $4$-antiprism. The corresponding links have homeomorphic complements, but the first link has $2$-fold branch covering space with hyperbolic structure, and the $2$-fold branch covering space contains incompressible tori corresponding to $4$-belts – cyclic sequences of $4$ facets such that two facets are adjacent if and only if they are successive (see more details in [E22]).
@@
Problem 2.
To characterise ideal right-angled $3$-polytopes corresponding to Hamiltonian cycles on compact right-angled $3$-polytopes.
@@
To characterise Hamiltonian cycles on simple $3$-polytopes corresponding to nonselfcrossing Eulerian cycles on ideal right-angled hyperbolic $3$-polytopes.
@@
To characterise Hamiltonian cycles on simple $3$-polytopes corresponding to nonselfcrossing Eulerian cycles on $4$-valent convex polytopes.
Problem 5.
@@
Problem 6.
To enumerate all combinatorially different nonselfcrossing Eulerian cycles on any ideal right-angled $3$-polytope.

### v2

@@ Abstract.
We consider Eulerian cycles without transversal selfintersections in $4$-valent planar graphs. We prove that any cycle of this type in the graph of an ideal right-angled hyperbolic $3$-polytope corresponds to a hyperbolic link such that its complement consists of $4$-copies of this polytope glued according to its checkerboard coloring. Moreover, this link consists of trivially embedded circles bijectively corresponding to the vertices of the polytope. For such cycles we prove that the $3$-antiprism $A(3)$ (octahedron) has exactly $2$ combinatorially different cycles, the $4$-antiprism $A(4)$ has exactly $7$ combinatorially different cycles, and these cycles correspond to $7$ cycles (perhaps combinatorially equivalent) on any polytope different from antiprisms, and any antiprism $A(k)$ has at least $2$ combinatorially different cycles. The $2$-fold branched covering space corresponding to our link is a small cover over some simple $3$-polytope. This small cover is defined by a Hamiltonian cycle on it. We show that any Hamiltonian cycle on a compact right-angled hyperbolic $3$-polytope arises in this way, while for a Hamiltonian cycle on a right-angled hyperbolic $3$-polytope of finite volume the necessary and sufficient condition is that at each ideal vertex it does not go straight. We introduce a transformation of a Eulerian cycle along conjugated vertices allowing to build new cycles from a given one.
@@ 1. Introduction
Our main result (Theorem 3.2) is that any Eulerian cycle without transversal selfintersections in the graph of any ideal right-angled $3$-polytope $P$ corresponds to a link whose circles bijectively correspond to vertices of $P$ and the complement is decomposed into $4$ copies of $P$. We show that the $2$-fold branched covering space corresponding to this link is a small cover build by the A.D. Mednykh [M90] construction from a Hamiltonian cycle on a simple $3$-polytope. We show that any Hamiltonian cycle on a compact right-angled hyperbolic $3$-polytope arises in this way, while for a Hamiltonian cycle on a right-angled hyperbolic $3$-polytope of finite volume the necessary and sufficient condition is that at each ideal vertex it does not go straight. We provide methods how to build Eulerian cycles without transversal selfintersections on any ideal right-angled $3$-polytope.
@@ 2. Basic facts
It was proved in [E22b] that the family of manifolds $\{M(P)\}$, where $M(P)$ is the double of the manifold obtained from $N(P)$ by adding the boundary torus at each cusp, is cohomologically rigid over $\mathbb{Z}_{2}$, that is two manifolds from this family are homeomorphic if and only if their cohomology rings over $\mathbb{Z}_{2}$ are isomorphic as graded rings.
@@
Construction 2.5 (A small cover and a link from a Hamiltonian cycle).
@@
If the components of $\partial Q\setminus\Gamma$ correspond to the pairs of colors $(1,2)$ and $(3,4)$, then $\tau=e_{1}+e_{2}=e_{3}+(e_{1}+e_{2}+e_{3})$ is the hyperelliptic involution on $N(Q,\widetilde{\Lambda}_{\Gamma})$. Moreover, the mapping $N(P,\widetilde{\Lambda}_{\Gamma})\to S^{3}$ is a $2$-fold branched covering with the following branch set (see details in [EE25, Section 4.5]). The edges of $Q$ not lying in $\Gamma$ form a perfect matching of the graph $G(Q)$ – a disjoint set of edges covering all the vertices. Then the preimage of this set in $N(Q,\widetilde{\Lambda}_{\Gamma})$ and in $S^{3}$ is a disjoint set of circles $C_{\Gamma}$, each circle glued of two copies of the corresponding edge. This link is the brach set of the covering. The detailed description of this link is given in [G24]. In [VM99S2] this construction was generalized from Hamiltonian cycles to Hamiltonian theta-graphs and Hamiltonian $K_{4}$-subgraphs.
@@ 4. Edge-twists and the existence of nonselfcrossing Eulerian cycles
5. Transformations of nonselfcrossing Eulerian cycles
Definition 5.1.
We will call two edges $E_{1}$ and $E_{2}$ of $Q$ not lying in a Hamiltonian cycle $\Gamma$ conjugated, if each edge intersects both components of the complement in $\Gamma$ to the vertices of the other edge (in other words, if $\Gamma\cup E_{1}\cup E_{2}$ is homeomorphic to the full graph $K_{4}$ on four vertices). We call two vertices of an ideal right-angled $3$-polytope $P$ conjugated along the nonselfcrossing Eulerian cycle $\gamma$, if the corresponding edges of $Q(P,\gamma)$ are conjugated.
@@ 5. Transformations of nonselfcrossing Eulerian cycles
The circles in $C_{\Gamma}$ corresponding to the edges of $Q$ not lying in $\Gamma$ are linked if and only if the edges are conjugated.
Proof.
This becomes evident if we look at the link in a way shown in Fig. 2 on the right. ∎
Lemma 5.3.
Each edge of $Q\setminus\Gamma$ has a conjugated edge.
Proof.
Indeed, let the edge $E$ of $Q\setminus\Gamma$ have no conjugated edges. $E$ is the intersection of two facets $F_{i}$ and $F_{j}$ of $Q$ lying in the closure of the same connected component of $\partial Q\setminus\Gamma$. Then both vertices of $E$ belong to the same facet $F_{k}$ lying in the closure of the other connected component. In this case $E$ belongs to $F_{k}$, which is a contradiction. ∎
Construction 5.4 (Transformation of a nonselfcrossing Eulerian cycle along conjugated vertices).
Given two conjugated vertices $v$ and $w$ of a nonselfcrossing Eulerian cycle $\gamma$ of an ideal right-angled $3$-polytope $P$ we can build a new nonselfcrossing Eulerian cycle in the following way. In both vertices we change the pairs of successive edges of the cycle to complementary pair (see Fig. 7). In Fig. 8 we show how the Hamiltonian cycle $\Gamma_{\gamma}$ is transformed under this operation (the new Hamiltonian cycle belongs to another polytope obtained from $Q$ by two flips).
Figure 7. Local transformation of the Eulerian cycle $\gamma$ and the corresponding flip of the polytope $Q$
Figure 8. Transformation of the Hamiltonian cycle $\Gamma_{\gamma}$.
Proposition 5.5.
Let $\gamma$ be a nonselfcrossing Eulerian cycle on the ideal right-angled $3$-polytope $P$. Then for any vertex of $P$ there is at least one conjugated vertex and the corresponding transformation of $\gamma$.
Conjecture 5.6.
Any two nonselfcrossing Eulerian cycles are connected by a sequence of transformations along conjugated vertices.
6. Links associated to Hamiltonian cycles on right-angled $3$-polytopes
A $k$-belt is a cyclic sequence of $k$ facets such that facets are adjacent if and only if they are successive and no three facets have a common vertex. It follows from results by A.V. Pogorelov and E.M. Andreev that a $3$-polytope is combinatorially equivalent to a compact right-angled hyperbolic $3$-polytope if and only if it is a simple polytope different from the simplex and has no $3$- and $4$-belts (see more details in [E19]). These polytopes are called Pogorelov polytopes.
Proposition 6.1.
Every compact right-angled hyperbolic $3$-polytope $Q$ with a Hamiltonian cycle $\Gamma$ defines an ideal right-angled polytope $P$ obtained by shrinking to points the edges of the perfect matching consisting of the edges not in the cycle. The polytope $Q$ has an induced nonselfcrossing Eulerian cycle $\gamma$ such that $\Gamma=\Gamma_{\gamma}$. For the link $C_{\gamma}$ both the complement $S^{3}\setminus C_{\gamma}$ and the $2$-fold branched covering space have complete hyperbolic structures obtained by gluing right-angled polytopes.
@@ 6. Links associated to Hamiltonian cycles on right-angled $3$-polytopes
The dodecahedron is a unique compact right-angled polytope with minimal number of facets (equal to $12$). Up to combinatorial symmetries it has a unique Hamiltonian cycle. In Fig. 9 we show this Hamiltonian cycle, and the way how the associated ideal right-angled polytope is obtained from $A(4)$ by a sequence of two restricted edge-twists. We also show another polytope corresponding to another Eulerian cycle on the $4$-antiprism. The corresponding links have homeomorphic complements, but the first link has the $2$-fold branched covering space with a hyperbolic structure, and the $2$-fold branched covering space of the second link contains incompressible tori corresponding to $4$-belts (see more details in [E22a]).
@@
Now consider right-angled hyperbolic $3$-polytopes of finite volume. Each finite vertex of such a polytope has valency $3$, and ideal vertices have valency $4$. Using Andreev’s theorem it can be shown that cutting off ideal vertices defines a bijection between combinatorial types of right-angled hyperbolic $3$-polytopes of finite volume and almost Pogorelov polytopes different from the $4$-prism (cube) and the $5$-prism (see [DO01, Theorem 10.3.1] and [E19, Theorem 6.5]). Moreover, all quadrangles of the resulting polytope arise from ideal vertices. A simple $3$-polytope is called an almost Pogorelov polytope, if it is different from the simplex, has no $3$-belts, and any its $4$-belt surrounds a quadrangular facet.
Proposition 6.5.
Let $\Gamma$ be a Hamiltonian cycle on an almost Pogorelov $3$-polytope $Q$. Then shrinking to points the edges of the perfect matching complementary to $\Gamma$ gives an ideal right-angled polytope $P$ if and only if each quadrangle of $Q$ has three edges in $\Gamma$. In this case the polytope $Q$ has an induced nonselfcrossing Eulerian cycle $\gamma$ such that $\Gamma=\Gamma_{\gamma}$. For the link $C_{\gamma}$ the complement $S^{3}\setminus C_{\gamma}$ is hyperbolic and the $2$-fold branched covering space becomes hyperbolic after cutting along incompressible Klein bottles corresponding to quadrangles of $Q$.
Proof.
For each quadrangle of $Q$ the Hamiltonian cycle contains either two its opposite edges, or three of its edges. In the first case after shrinking the quadrangle becomes a bigon and we obtain no polytope. If each quadrangle intersects $\Gamma$ by three edges, then it intersects at vertices two edges of the complementary perfect matching, and by  [E19, Corollary 12.31] cutting off all the edges of the matching gives an almost Pogorelov polytope different from the $4$- and the $5$-prism and produces all its quadrangles. Then shrinking quadrangles to points we obtain an ideal right-angled hyperbolic $3$-polytope $P$. The same polytope we obtain by shrinking all the edges of the matching. By [E22a, Theorem 4.12] quadrangles of $Q$ correspond to incompressible Klein bottles in $N(Q,\widetilde{\Lambda}_{\Gamma})$ such that the complement to their union has a complete hyperbolic structure of finite volume. ∎
Corollary 6.6.
A Hamiltonian cycle on a right-angled hyperbolic $3$-polytope of finite volume corresponds to a nonselfcrossing Eulerian cycle on the ideal right-angled hyperbolic $3$-polytope if and only if at each ideal vertex it turns left or right, but does not go straight.
Example 6.7.
In Fig. 10 we show a Hamiltonian cycle on the $3$-dimensional permutohedron intersecting each quadrangle by $3$ edges. After shrinking quadrangles to points we obtain a Hamiltonian cycle on the ideal octahedron.
Figure 10. The Hamiltonian cycle on the permutohedron corresponding to the Hamiltonian cycle on the ideal octahedron.
Problem 2.
To characterise ideal right-angled $3$-polytopes corresponding to Hamiltonian cycles on (a) compact right-angled hyperblolic $3$-polytopes (b) right-angled hyperblolic $3$-polytopes of finite volume.
Remark 6.8.
In [E19, Theorem 9.17] is was proved that any ideal right-angled hyperbolic $3$-polytope $P$ can be obtained from an almost Pogorelov polytope or the polytope $P_{8}$ (drawn in the center at the bottom of Fig. 2) by a contraction of edges of a perfect matching such that no quadrangle contains two edges of the matching. Nevertheless, the complement to this matching may be not a Hamiltonian cycle, but a union of cycles containing all the vertices of the polytope.
@@
To characterise Hamiltonian cycles on simple $3$-polytopes corresponding to nonselfcrossing Eulerian cycles on (a) ideal right-angled hyperbolic $3$-polytopes (b) $4$-valent convex polytopes.